# Sequential model confidence sets

### v3

@@ Appendix D SMCSs by FDR control
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$. Following Wang and Ramdas (2022), we let $E_{[1],t}\geq\dots\geq E_{[m],t}$ be the order statistics of $E_{1,t},\dots,E_{m,t}$ in decreasing order and define $i^{\star}_{t}=\max\{i\in\{1,\dots,m\}\;|\;{iE_{[i],t}}\geq m/{\alpha}\}$, for a given $\alpha\in(0,1)$. Then, we have a time-uniform FDR control at level $\alpha$ for the sequential tests $\Psi_{i,t}=(\mathbbm{1}\{E_{i,t}\geq E_{[i_{t}^{\star}],t}\}),i=1,\dots,m,t\in\mathbb{N}.$ As our construction for the strong and uniformly weak hypothesis builds on e-processes, we can use the above sequential tests with respect to the merged e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ given in Section 3.2. That is, at each $t\in\mathbb{N}$, we reject the $i^{\star}_{t}$ hypotheses with the largest values of the corresponding e-processes, and obtain the alternative SMCS ˇ$\check{\mathcal{M}}_{t}=\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<E_{[i_{t}^{\star}]\cdot,t}\}$. Importantly, this sequence does not necessarily satisfy the time-uniform coverage property from Section 2 anymore. In contrast, it bounds the expected ratio of wrongly rejected models to the total number of rejected models at any random time by $\alpha$, that is, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$ and any random stopping time $\tau$, we have $\mathbb{E}_{\mathbb{Q}}\big{[}{\lvert\check{\mathcal{M}}^{c}_{\tau}\cap\mathcal{M}^{\star}_{\tau}\rvert/\max\{1,\lvert\check{\mathcal{M}}_{\tau}^{c}\rvert\}}\big{]}\leq\alpha,$ where ˇ$\check{\mathcal{M}}^{c}_{\tau}=\{1,\dots,m\}\setminus\check{\mathcal{M}}_{\tau}$ and $(\mathcal{M}^{\star})_{t\in\mathbb{N}}$ is given by (1) or (2).
Appendix E Proofs for Section 3.3
@@
Appendix F Further simulation results
F.1 Uniformly weakly superior objects for unbounded losses
@@
F.2 Strongly superior objects for non-Gaussian data
@@
    If $H_{0,\widehat{\mathcal{M}}_{t-1}}$ is not rejected, then we set ^$\widehat{\mathcal{M}}_{t}=\widehat{\mathcal{M}}_{t-1}$ and $t\leftarrow t+1$, and we go back to step 1. Else, we eliminate the worst-performing model, namely, the one indexed by
@@
    We then set ^$\widehat{\mathcal{M}}_{t-1}\leftarrow\widehat{\mathcal{M}}_{t-1}\backslash\{e_{\text{R,M}}\}$ and go back to step 2.
@@
In the above implementation of the MCS algorithm, after a model is eliminated, it is totally discarded in future time points. That is, when we set $t\leftarrow t+1$ and move back to step 1, we only repeat the process for the models in ^$\widehat{\mathcal{M}}_{t}$, and not for all models in $\mathcal{M}_{0}$. For anytime-valid methods, this is not possible, as illustrated by Equation (6), where $m$ stays constant over time.

### v4

@@ Appendix E SMCSs by FDR control
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$. Following Wang and Ramdas (2022), we let $E_{[1],t}\geq\dots\geq E_{[m],t}$ be the order statistics of $E_{1,t},\dots,E_{m,t}$ in decreasing order and define $i^{\star}_{t}=\max\{i\in\{1,\dots,m\}\;|\;{iE_{[i],t}}\geq m/{\alpha}\}$, for a given $\alpha\in(0,1)$. Then, we have a time-uniform FDR control at level $\alpha$ for the sequential tests $\Psi_{i,t}=(\mathbbm{1}\{E_{i,t}\geq E_{[i_{t}^{\star}],t}\}),i=1,\dots,m,t\in\mathbb{N}.$ As our construction for the strong and uniformly weak hypothesis builds on e-processes, we can use the above sequential tests with respect to the merged e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ given in Section 3.2. That is, at each $t\in\mathbb{N}$, we reject the $i^{\star}_{t}$ hypotheses with the largest values of the corresponding e-processes, and obtain the alternative SMCS $\check{\mathcal{M}}_{t}=\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<E_{[i_{t}^{\star}]\cdot,t}\}$. Importantly, this sequence does not necessarily satisfy the time-uniform coverage property from Section 2 anymore. In contrast, it bounds the expected ratio of wrongly rejected models to the total number of rejected models at any random time by $\alpha$, that is, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$ and any random stopping time $\tau$, we have $\mathbb{E}_{\mathbb{Q}}\big[{\lvert\check{\mathcal{M}}^{c}_{\tau}\cap\mathcal{M}^{\star}_{\tau}\rvert/\max\{1,\lvert\check{\mathcal{M}}_{\tau}^{c}\rvert\}}\big]\leq\alpha,$ where $\check{\mathcal{M}}^{c}_{\tau}=\{1,\dots,m\}\setminus\check{\mathcal{M}}_{\tau}$ and $(\mathcal{M}^{\star})_{t\in\mathbb{N}}$ is given by (1) or (2).
Appendix F Proofs for Section 3.3
@@
Appendix G Further simulation results
G.1 Uniformly weakly superior objects for unbounded losses
@@
G.2 Strongly superior objects for non-Gaussian data
@@
    If $H_{0,\widehat{\mathcal{M}}_{t-1}}$ is not rejected, then we set $\widehat{\mathcal{M}}_{t}=\widehat{\mathcal{M}}_{t-1}$ and $t\leftarrow t+1$, and we go back to step 1. Else, we eliminate the worst-performing model, namely, the one indexed by
@@
[…]
@@
In the above implementation of the MCS algorithm, after a model is eliminated, it is totally discarded in future time points. That is, when we set $t\leftarrow t+1$ and move back to step 1, we only repeat the process for the models in $\widehat{\mathcal{M}}_{t}$, and not for all models in $\mathcal{M}_{0}$. For anytime-valid methods, this is not possible, as illustrated by Equation (6), where $m$ stays constant over time.
@@
Appendix H Accelerated algorithm for e-value adjustment
In this section, we present an accelerated version of the e-value adjustment algorithm of Vovk and Wang (2021). The idea and its rigorous formulation is due to Tim Stephan, a student seminar participant at ETH Zürich.
The complexity of the original algorithm (Vovk and Wang, 2021, Algorithm 1) is $O(m^{2})$, where $m$ is the number of e-values used as an input. A quick look at this algorithm (Algorithm 1) reveals that the quadratic complexity comes from the nested loop, which runs a total of $m^{2}$ iterations.
Observe that the inner loop does not need to run all the way from $1$ to $m$. Due to the sorting step, the minimum term is attained for some value $i\in\{0,\ldots,k-1\}$. For larger indices, the terms that are added to the numerator are larger than $e_{\pi(k)}$, so the overall fraction takes higher values. This observation can reduce the number of iterations to $1+\ldots+m-1$, which is however still quadratic in $m$.
Algorithm 1 E-Value Adjustment Algorithm (Vovk and Wang, 2021)
1:A sequence of e-values $e_{1},\ldots,e_{m}$.
2:Find a permutation $\pi:\{1,\ldots,m\}\to\{1,\ldots,m\}$ such that $e_{\pi(1)}\leq\ldots e_{\pi(m)}$.
3:Set $S_{0}\vcentcolon=0$.
4:for $i=1,\ldots,m$ do
5:  $S_{i}\vcentcolon=S_{i-1}+e_{\pi(i)}$.
6:end for
7:for $k=1,\ldots,m$ do
8:  $e_{\pi(k)}^{\star}\vcentcolon=e_{\pi(k)}$.
9:  for $i=1,\ldots,m$ do
10:   $e\vcentcolon=\frac{e_{\pi(k)}+S_{i}}{i+1}$
11:   if $e<e_{\pi(k)}^{\star}$ then
12:     $e_{\pi(k)}^{\star}\vcentcolon=e$.
13:   end if
14:  end for
15:end for
16:Adjusted e-values $e_{1}^{\star},\ldots,e_{m}^{\star}$.
It turns out that we can actually take advantage of the fact that the e-values are first being sorted to bring the number of iterations down to a linear level. In that case, the dominating part will be the ordering of the e-values, whose complexity is generally $O(m\log m)$. The accelerated e-value adjustment algorithm relies on the following lemma:
Lemma H.1.
Let $a_{1}\leq\ldots\leq a_{m}$ be positive real numbers. Set $S_{0}=0$ and $S_{i}=\sum_{k=1}^{i}a_{k}$ for $i=1,\ldots,m$. For all $k\in\{1,\ldots,m\}$, define
$Q_{k,i}=\frac{a_{k}+S_{i}}{i+1},\quad i=0,\ldots,k-1.$
Then:
  1. 1.
    For each $k\in\{1,\ldots,m\}$, there exists an index $i_{k}\in\{1,\ldots,k-1\}$ such that $Q_{k,i-1}\geq Q_{k,i}$ for all $i\leq i_{k}$ and $Q_{k,i}<Q_{k,i+1}$ for $i\geq i_{k}$. By convention, we choose $i_{k}$ to be the largest index with this property.
  2. 2.
    For all $k\in\{1,\ldots,m-1\}$, it holds that $i_{k}\leq i_{k+1}$.
Proof.
The proof is based on the two following straightforward claims. Let $x_{1},\ldots,x_{m},y$ be arbitrary real numbers and let $\overline{x}_{1:m}$ denote the arithmetic mean of $x_{1},\ldots,x_{m}$. Then,
  • If $y\geq\overline{x}_{1:m}$, then
    $y\geq\frac{x_{1}+\ldots+x_{m}+y}{m+1}\geq\overline{x}_{1:m},$
    with equality if and only if $y=\overline{x}_{1:m}$.
  • If $y\leq\overline{x}_{1:m}$, then
    $y\leq\frac{x_{1}+\ldots+x_{m}+y}{m+1}\leq\overline{x}_{1:m},$
    with equality iff $y=\overline{x}_{1:m}$.
We now prove the two parts of the lemma separately.
  1. 1.
    Starting with $Q_{k,1}$, and since $a_{i}\leq a_{i+1}$ for all $i\in\{1,\ldots,m-1\}$, the sequence $Q_{k,1},Q_{k,2},\ldots$ is decreasing until we hit an index $i\in\{1,\ldots,k-1\}$ such that $a_{i+1}>Q_{k,i}$. We call this index $i_{k}$. If no such index exists, we set $i_{k}=k-1$. We now need to show that the sequence $Q_{k,i},Q_{k,i+1},\ldots$ is strictly increasing for $i\geq i_{k}$. If $i_{k}=k-1$, there is nothing to show because $Q_{k,i}$ is not defined for $i\geq k$. If $i_{k}<k-1$, then $a_{i_{k}+1}>Q_{k,i_{k}}$, yields that $a_{i_{k}+1}>Q_{k,i_{k}+1}$. The inequality $Q_{k,i_{k}}<a_{i_{k}+1}$ shows that $Q_{k,i_{k}}<Q_{k,i_{k}+1}<a_{i_{k}+1}\leq a_{i_{k}+2}$. In turn, this yields $Q_{k,i_{k}+1}<Q_{k,i_{k}+2}<a_{i_{k}+2}\leq a_{i_{k}+3}$, so the claim follows by induction.
  2. 2.
    Fix $k\in\{1,\ldots,m-1\}$. If $i_{k}=1$, there is nothing to show. By definition, it holds that $Q_{k,i_{k}-1}\geq Q_{k,i_{k}}$. This implies that $a_{i_{k}}\leq Q_{k,i_{k}-1}$ (this follows by the two facts in the beginning of the proof). On the other hand, it holds that $a_{j}\leq a_{j+1}$ for all indices $j\in[k-1]$, so $a_{i_{k}}\leq Q_{k,i_{k}-1}\leq Q_{k+1,i_{k}-1}$. This yields $Q_{k+1,i_{k}-1}\geq Q_{k+1,i_{k}}$, so the turning point of the sequence $Q_{k+1,1},Q_{k+1,2},\ldots$ is at least as large as $i_{k}$.
This lemma reveals that the inner loop needs to run for even fewer iterations. Once we reach the value $i_{k}$, we can update the value of $k$ and start the index $i$ from $i_{k}$ instead of starting from $1$. We will update the value of $k$ again when we reach the point where the sequence $Q_{k,i}$ becomes strictly increasing. This yields a total of $(i_{k}+1)-(i_{k-1}-1)$ iterations for each value of $k\in\{2,\ldots,m\}$ and a total of $i_{1}+1$ iterations for $k=1$. Therefore, the total number of iterations of the nested loop is
$(i_{1}+1)+(i_{2}+1-i_{1}+1)+\ldots(i_{m}+1-i_{m-1}+1)=i_{m}+2m-1<3m,$
which is linear in $m$. This idea is summarized in Algorithm 2, which is an accelerated version of Algorithm 1.
Algorithm 2 Accelerated E-Value Calibration
1:A sequence of e-values $e_{1},\ldots,e_{m}$.
2:Find a permutation $\pi:\{1,\ldots,m\}\to\{1,\ldots,m\}$ such that $e_{\pi(1)}\leq\ldots e_{\pi(m)}$.
3:Set $S_{0}\vcentcolon=0$.
4:for $i=1,\ldots,m$ do
5:  $S_{i}\vcentcolon=S_{i-1}+e_{\pi(i)}$.
6:end for
[…]
9:for $2\leq k\leq m$ do
10:  $e_{\pi(k)}^{\star}\leftarrow\frac{e_{\pi(k)}+S_{i}}{i+1}$.
11:  while i<k1i<k-1 do
12:   $i\leftarrow i+1$.
13:   $e_{\text{temp}}\leftarrow\frac{e_{\pi(k)}+S_{i}}{i+1}$.
14:   if $e_{\text{temp}}\leq e_{\pi(k)}^{\star}$ then
15:     $e_{\pi(k)}^{\star}\leftarrow e_{\text{temp}}$.
16:   else
17:     $i\leftarrow i-1$.
18:     break
19:   end if
20:  end while
21:end for
22:Adjusted values $e_{1}^{\star},\ldots,e_{m}^{\star}$.